OpenStudy (anonymous):

"Suppose you receive an inheritance of $28,000. You decide to invest in stocks and bonds. To diversify, you decide that three times the amount in bonds should equal four times the amount in stocks. How much should be invested in bonds? How much should be invested in stocks?"

OpenStudy (anonymous):

OK. Simple investment diversification problem. Here's what we know from the word problem. Let S = $ amount in Stocks Let B = $ amount in Bonds $28,000 = total investment S + B = $28,000 3B = 4S Solve for B 3B/4 + B = 28,000 3B + 4B = 112000 - Clearing fractions 7B = 112000 B= 16000 Solve for S B + S = 28,000 16000 + S = 28000 - Substitute in the B S = 12,000 Check your work 3B = 4S ? 3(16000) = 4(12000) ? 48,000 = 48,000 ? YES $16,000 will be invested in Bonds, while $12,000 will be invested in stocks.

OpenStudy (anonymous):

youre a lifesaver! thank you!